3(2x + 6) - 5(x + 8) = x - 22

I really really would like to know how to solve this I have multiple questions on as assesment in a class I just started taking and i just do not remember how to solve this so any sort of help would be really appreciated. Oh , and i'm not getting graded seriously for it I just reall would like to know, we're only getting graded for doing it, but we will be taing a test on it in the future so i'm curious now. Thank you kindly!

Answer by Gogonati(855)

3(2x+6)-5(x+8)=x-22
6x+18-5x-40=x-22
x-22=x-22 <=> 0*x=0 The last equality is an identity, this means is an expression that satisfies for each value of x. Therefore this equation has infinity solutions or x={x| x∈R} the set of real numbers, because is a polynomial.